# Stochastically Structured Reservoir Computers for Financial and Economic System Identification

### v1

@@ Abstract
This paper introduces a methodology for identifying and simulating financial and economic systems using stochastically structured reservoir computers (SSRCs). The proposed framework leverages structure-preserving embeddings and graph-informed coupling matrices to model inter-agent dynamics with enhanced interpretability. A constrained optimization scheme ensures that the learned models satisfy both stochastic and structural constraints. Two empirical case studies, a dynamic behavioral model of resource competition among agents, and regional inflation network dynamics, illustrate the effectiveness of the approach in capturing and anticipating complex nonlinear patterns and enabling interpretable predictive analysis under uncertainty.
@@ 1 Introduction
In this article, we propose a methodology that introduces stochastically structured reservoir computers that are tailored to identify and simulate financial and economic systems. Central to this approach is the construction of a reservoir computing framework in which the architecture of the output coupling matrix is informed by observed interactions among dynamic agents within the system. This structure is represented through a relational graph that captures these inter-agent dynamics. By embedding the original signals into a higher-dimensional space through stochastic dynamic embeddings, the structure-preserving reservoir computing model improves the system identification process. The training phase involves solving a constrained optimization problem to estimate the output coupling matrix, ensuring it adheres to the predefined structural constraints. This design not only preserves the stochastic properties of the system, but also enables a more interpretable and efficient modeling of complex financial and economic behaviors.
The proposed methodology is applied through two empirical case studies: (i) the evolution of resource competition behavior among agents, and (ii) the inflation network dynamics among countries in the CAPARD region (Central America, Panama, and the Dominican Republic), the United States, and China. The latter case examines the contribution of US conventional monetary policy, via changes in the federal funds rate, to domestic and regional inflation, within a time frame that includes the COVID-19 pandemic and its economic aftermath. By considering stochastic and structural constraints, the methodology connects theoretical expressiveness with practical interpretability, allowing the model to reflect both the uncertainty and the internal logic of the systems it represents.
In light of recent advances in reservoir computing, particularly those discussed in “Emerging Opportunities and Challenges for the Future of Reservoir Computing”, this study makes several key contributions. First, it strengthens the mathematical foundation of reservoir computing by integrating stochastic structures, thereby enhancing the framework’s capability to model complex economic and financial systems. Second, the methodology provides a practical algorithmic solution that balances computational efficiency with modeling accuracy, resonating with the pursuit of lightweight and fast-adapting models suitable for real-time economic forecasting. Lastly, by focusing on financial applications, this research extends the applicability of reservoir computing to domains where data-driven decision making is critical, thus bridging the gap between theoretical advances and industrial adoption.
The structure of the paper is as follows: we begin by establishing the notation and preliminaries, followed by the development of the proposed methodology. We then present two empirical case studies that illustrate the effectiveness of our approach, and conclude with a discussion of the results and future directions.
@@ 2 Preliminaries and Notation
The inverse, $\operatorname{vec}^{\dagger}(c)$, for $c\in\mathbb{R}^{n^{2}}$, reconstructs $X$ with:
$x_{i,j}=c_{(j-1)n+i},\quad i,j=1,\dots,n.$
Thus, $\operatorname{vec}^{\dagger}(\operatorname{vec}(X))=X$.
@@
[…]
represents the class of stochastic (column-stochastic) matrices in $\mathbb{R}^{m\times n}$.The symbol $e_{j,k}(n)$ will denote the matrix in $\mathbb{R}^{n\times n}$ with $j,k$ entry equal to one and with zeros elsewhere.
@@ 3 Methodology
The dynamic models considered for this study are determined by generic switched control systems of the form:
$\mathbf{x}(t+1)$ $=W_{\tau(t)}\eth_{p}\left(\begin{bmatrix}\mathbf{x}(t)\\ \mathbf{u}(t)\end{bmatrix}\right)+\mathbf{e}_{t},$ (1)
$\mathbf{y}(t)$ $=C_{\tau(t)}\mathbf{x}(t)+\mathbf{r}_{t}$ (2)
with $\mathbf{e}_{t}\sim N(\mathbf{0},\Sigma_{e})$ and $\mathbf{r}_{t}\sim N(\mathbf{0},\Sigma_{r})$.
The models (1), (2) correspond to what we call in this study a stochastically structured regressive reservoir computers whose architecture is determined by a generalization of the models presented in Vides et al. (2025).
Let us write ð~$\tilde{\eth}_{p}(x)$ to denote the map ð$\eth_{p}:\mathbb{R}^{n}\to\mathbb{R}^{d_{p}(n)}$ for $d_{p}(n)=n(n^{p}-1)/(n-1)+1$, that is determined by the following expression:
ð~$\tilde{\eth}_{p}(x):=\frac{1}{p+1}\begin{bmatrix}x^{\otimes 1}\\ x^{\otimes 2}\\ \vdots\\ x^{\otimes p}\\ 1\end{bmatrix}$ (3)
Here, the number $p$ will be called the order of the embedding map ð~$\tilde{\eth}_{p}$.
@@
The embedding map ð~$\tilde{\eth}_{p}(x)$ preserves stochastic vectors.
@@
It can be seen that for any $x=[x_{j}]\in\mathbb{S}_{n,1}(\mathbb{R})$, if we define $y=x\otimes x$, then $y\in\mathbb{R}^{n^{2}}$ and we will have that:
$\mathbf{1}_{n^{2}}^{\top}y$ $=\sum_{j=1}^{n}x_{j}\sum_{k=1}^{n}x_{k}=\sum_{j=1}^{n}x_{j}\cdot 1=\sum_{j=1}^{n}x_{j}=1.$
This implies that for any $p$, $x^{\otimes p}$ is stochastic, and by the definition of ð~$\tilde{\eth}_{p}$, we will have that the sum of the entries of ð$\eth_{p}(x)$ equals $(p+1)/(p+1)=1$. Consequently, the sum of the entries of ð~$\tilde{\eth}_{p}(x)$ equals 1. This completes the proof.
From here on, the map ð~$\tilde{\eth}_{p}$ will be called a stochastic $p$-embedding.
@@
Given positive integers $n,p$. There are an integer $0<\rho_{p}(n)<d_{p}(n)$ and a sparse matrix $R_{p}(n)\in\mathbb{R}^{\rho_{p}(n)\times d_{p}(n)}$ with $d_{p}(n)$ nonzero entries, such that $R_{p}(n)\>\tilde{\eth}_{p}(x)$ is stochastic and has the least number of non-redundant words (monomial terms) for any $x\in\mathbb{R}^{n}$.
@@
Let $n,p$ be positive integers. Consider the structured embedding map ð~$\tilde{\eth}_{p}:\mathbb{R}^{n}\to\mathbb{R}^{d_{p}(n)}$, where $d_{p}(n)=n(n^{p}-1)/(n-1)+1$ corresponds to the total number of distinct tensor monomials up to degree $p$, plus a constant term.
We aim to construct a sparse matrix $R_{p}(n)\in\mathbb{R}^{\rho_{p}(n)\times d_{p}(n)}$ that maps the embedding ð~$\tilde{\eth}_{p}(x)$ to a stochastic vector of reduced dimension, while preserving all non-redundant monomial terms.
Let us start by defining the matrix $R\in\mathbb{R}^{1\times d_{p}(n)}$, with $1$ in its $R_{11}$ entry and with all other entries equal to zero.
For each $2\leq j\leq d_{p}(n)$, let us consider the indices $j=k_{1}(j)<k_{2}(j)<\cdots<k_{n_{j}}(j)<d_{p}(n)$ that correspond to the same monomial in ð~$\tilde{\eth}_{p}(x)$, let us define the matrix $R_{0}\in\mathbb{R}^{1\times d_{p}(n)}$ with $1$ in its $R_{1k_{l}(j)}$ entries for $1\leq l\leq n_{j}$, and with all other entries equal to zero. Let us now define the augmented matrix
@@
where $R^{\prime}\in\mathbb{R}^{1\times d_{p}(n)}$ is the matrix with entry $R^{\prime}_{1d_{p}(n)}$ equal to $1$, and with all other entries equal to 00.
Let us set $R_{p}(n):=R$. It can be seen by the way $R$ has been constructed, that the operation $R_{p}(n)\tilde{\eth}_{p}(x)$ assigns each group of duplicates in ð~$\tilde{\eth}_{p}(x)$ to a single representative coordinate of $R_{p}(n)\tilde{\eth}_{p}(x)$. This selection is performed by adding over the redundant entries and projecting onto a reduced subspace. Because of this, and as a consequence of Lemma 1, it is clear that $R_{p}(n)\tilde{\eth}_{p}(x)$ is stochastic. Let us set $\rho_{p}(n)$ as the number of rows of $R$. This completes the proof.
Using the matrix $R_{p}(n)$ described in Lemma 2, we define the reduced stochastic embedding ð~$\tilde{\eth}_{p,r}$ as
ð~$\tilde{\eth}_{p,r}(x):=R_{p}(n)\tilde{\eth}_{p}(x),$
@@
For formal implementation purposes, the operation $R_{p}(n)\tilde{\eth}_{p}(x)$ can be equivalently represented by a vector composed of the distinct monomials of degree less than or equal to $p$, each multiplied by the appropriate integer scaling factor that accounts for its multiplicity in the original tensor expansion.
In this context, by a structured reservoir computer we mean a reservoir computer like the ones described in (1) and Vides et al. (2025) with ð$\eth_{p}:=\tilde{\eth}_{p,r}$, and whose output coupling matrix $W$ (in the sense of (Vides et al., 2025, (III.14))) is stochastic and has a structure that is determined by some relational graph $\mathcal{G}_{\mathcal{S}}=(V_{\mathcal{S}},E_{\mathcal{S}})$, representing the observed interaction between dynamic agents involved in the financial process under consideration, from $t$ to $t+1$ for each time-step in a given training time frame $\{0,1,\ldots,T\}$ under consideration, for some $T>0$.
@@
$\mathbb{B}_{\mathcal{S}}(n):=\left\{e_{j,k}(n)\left|(j,k)\in E_{\mathcal{S}}\right.\right\},$ (4)
then the existence of a $\mathbb{B}_{\mathcal{S}}(n)$-structured stochastic output coupling matrix $W$ that satisfies (1), is a consequence of the following theorem.
@@
The coupling matrices $W_{\tau(t)}$ in equation (1) can be approximately identified in the matrix set $\left(\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(n)\right)\cap\mathbb{S}_{m,n}(\mathbb{R})$.
@@
Let $\{\mathbf{x}(t)\in\mathbb{R}^{n}\mid t=0,1,\ldots,T\}$ denote a vector time series corresponding to the system’s evolution. We begin by applying a reduced embedding transformation ð~$\tilde{\eth}_{p,r}:=R_{p}(n)\circ\tilde{\eth}_{p}$ to the observed states, where $p$ is a prescribed tensor order and $R_{p}(n)$ is the structure-preserving compression map guaranteed by Lemma 2. This yields a compressed representation of each state, which we collect into the input data matrix
$\mathbf{X}_{0}(T):=\begin{bmatrix}|&&|\\ \tilde{\eth}_{p,r}(\mathbf{x}(0))&\cdots&\tilde{\eth}_{p,r}(\mathbf{x}(T-1))\\ |&&|\end{bmatrix},$
while the corresponding time-shifted outputs are stored in
$\mathbf{X}_{1}(T):=\begin{bmatrix}|&&|\\ \mathbf{x}(1)&\cdots&\mathbf{x}(T)\\ |&&|\end{bmatrix}.$
To ensure that the identified coupling matrix respects the structural constraints of the system, we consider a structured dictionary $\mathbb{B}_{\mathcal{S}}(n)=\{S_{1},\ldots,S_{q}\}\subset\mathbb{R}^{m\times n}$, derived from the graph $\mathcal{S}$ that encodes allowable interactions. Each $S_{j}$ acts as a basis component capturing a localized or interpretable mode of coupling. We construct a design matrix $X_{0}$ whose columns correspond to the vectorized actions of these basis elements on the embedded data:
$X_{0}:=\begin{bmatrix}\text{vec}(S_{1}\mathbf{X}_{0}(T))&\cdots&\text{vec}(S_{q}\mathbf{X}_{0}(T))\end{bmatrix},$
and define the target vector as $X_{1}:=\text{vec}(\mathbf{X}_{1}(T))$.
The identification task is then posed as the following structured matrix equation:
$\begin{bmatrix}X_{0}^{\top}X_{0}\\ C\end{bmatrix}\mathbf{a}=\begin{bmatrix}X_{0}^{\top}X_{1}\\ \mathbf{1}_{p}\end{bmatrix},$
where $C$ is a constraint matrix chosen to enforce the membership of the resulting linear combination $\hat{W}:=\sum_{j=1}^{q}\mathbf{a}[j]S_{j}$ in the structured subset $\mathbb{S}_{m,n}(\mathbb{R})$. Such constraints may include nonnegativity, block sparsity, or stochasticity properties, depending on the system’s prior assumptions.
As discussed in Boutsidis and Drineas (2009), the above system corresponds to a convex quadratic optimization problem with linear constraints (e.g., a Nonnegative Least Squares problem), and is therefore solvable in polynomial time up to arbitrary precision.
Since $\hat{W}$ is constructed as a linear combination of elements in $\mathbb{B}_{\mathcal{S}}(n)$, it lies in $\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(n)$ by definition. The enforcement of structural constraints via $C$ ensures that $\hat{W}\in\mathbb{S}_{m,n}(\mathbb{R})$. Hence,
$\hat{W}\in\left(\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(n)\right)\cap\mathbb{S}_{m,n}(\mathbb{R}),$

### v3

@@ Abstract
This paper introduces a methodology for identifying and simulating financial and economic systems using stochastically structured reservoir computers (SSRCs). The framework combines structure-preserving embeddings with graph-informed coupling matrices to model inter-agent dynamics while enhancing interpretability. A constrained optimization scheme guarantees compliance with both stochastic and structural constraints. Two empirical case studies, a nonlinear stochastic dynamic model and regional inflation network dynamics, demonstrate the effectiveness of the approach in capturing complex nonlinear patterns and enabling interpretable predictive analysis under uncertainty.
@@ 1 Introduction
This paper introduces a methodology for identifying and simulating financial and economic systems using stochastically structured reservoir computers (SSRCs). The approach builds on reservoir computing while constraining the output coupling matrices through relational graphs that reflect observed inter-agent interactions. Combined with stochastic embeddings, this design preserves structural properties and improves interpretability in system identification.
The methodology is illustrated through two empirical case studies: (i) a nonlinear stochastic dynamic model that highlights how SSRCs capture probabilistic dynamics and feedback in adaptive systems, and (ii) regional inflation dynamics across the CAPARD region (Central America, Panama, and the Dominican Republic), the United States, and China. The latter highlights the role of U.S. monetary policy, via changes in the federal funds rate, in shaping domestic and regional inflation during and after the COVID-19 pandemic.
Relative to recent advances in reservoir computing (Yan et al., 2024; Ehlers et al., 2025), this work makes three contributions. First, it integrates stochastic matrix structures into the framework, enhancing its capacity to capture nonlinear and uncertain economic behaviors. Second, it develops a constrained optimization scheme that balances structural compliance with computational efficiency. Third, by focusing on financial and economic applications, it extends the scope of reservoir computing to domains where interpretable, data-driven modeling is essential for policy and decision making.
The remainder of the paper is organized as follows. Section §2 introduces the preliminaries and notation. Section §3 presents the proposed methodology, followed by algorithms and interpretive discussion in Section §4. Section §5 reports simulation results for the two case studies, and Section §6 concludes with future directions.
@@ 2 Preliminaries and Notation
$\mathbb{S}_{m,n}(\mathbb{R}):=\left\{A\in(\mathbb{R}^{+}_{0})^{m\times n}\ \middle|\ \mathbf{1}_{m}^{\top}A=\mathbf{1}_{n}^{\top}\right\}$
represents the class of stochastic (column-stochastic) matrices in $\mathbb{R}^{m\times n}$. We will say that a vector $x\in(\mathbb{R}_{0}^{+})^{n}$ is substochastic if $\mathbf{1}_{n}^{\top}x\leq 1$. The symbol $e_{j,k}(m,n)$ will denote the matrix in $\mathbb{R}^{m\times n}$ with $j,k$ entry equal to one and with zeros elsewhere.
@@ 3 Methodology
The dynamic models considered in this study are described by discrete-time systems of the form
$y(t)$ $=W\,\eth_{p}(x(t))+e_{t},$ (1)
where, for each $t>0$, the vectors $x(t)$ and $y(t)$ are stochastic vectors of appropriate dimensions, and the term $e_{t}:=y(t)-W\,\eth_{p}(x(t))$ denotes the residual component.
The models in (1) correspond to what we refer to in this study as stochastically structured reservoir computers. Their architecture generalizes the framework introduced in modelos_reservorios, with $W$ constrained to be a column-stochastic matrix and with the mapping $\eth_{p}$ satisfying the property that $\eth_{p}(z)$ is stochastic whenever $z$ is a stochastic vector of appropriate dimension.
For this study we focus on a specific family of mappings $\eth_{p}$. Given a vector $s=[s_{1}~\cdots~s_{p+1}]^{\top}\in\{0,1\}^{p+1}$, let $\tilde{\eth}_{s,p}(x):\mathbb{R}^{n}\to\mathbb{R}^{d_{p}(n)}$ denote the map
$\tilde{\eth}_{s,p}(x):=\frac{1}{m}\begin{bmatrix}s_{1}\,x\\ s_{2}\,x^{\otimes 2}\\ \vdots\\ s_{p}\,x^{\otimes p}\\ s_{p+1}\end{bmatrix}$,
where $d_{p}(n)=\frac{n(n^{p}-1)}{n-1}+1$, and $m=\sum_{i=1}^{p+1}s_{i}$ denotes the number of active blocks. The parameter $m$, defined earlier, denotes the number of block terms considered. The selection of terms may be guided by exploratory analysis of prior signals or by domain knowledge. The integer $p$ will be referred to as the maximum order of the embedding map $\tilde{\eth}_{s,p}$. From here on, and for computational purposes, $\tilde{\eth}_{s,p}(x)$ will denote the expression obtained after eliminating the zero blocks from its formal definition. Under this consideration, the resulting map $\tilde{\eth}_{s,p}$ can now be seen as a mapping $\tilde{\eth}_{s,p}:\mathbb{R}^{n}\to\mathbb{R}^{d_{s,p}(n)}$, for $d_{s,p}(n):=s_{p+1}+\sum_{k=1}^{p}s_{k}n^{k}$.
@@
The embedding map $\tilde{\eth}_{s,p}$ preserves stochastic vectors.
@@
It can be seen that for any $x,y\in\mathbb{S}_{n,1}(\mathbb{R})$, if we define $z=x\otimes y$, then $z\in\mathbb{R}^{n^{2}}$ and we will have that:
$\mathbf{1}_{n^{2}}^{\top}z$ $=\sum_{j=1}^{n}x[j]\sum_{k=1}^{n}y[k]=\sum_{j=1}^{n}x[j]\cdot 1=\sum_{j=1}^{n}x[j]=1.$
This implies that for any $p$, $x^{\otimes p}$ is stochastic, and by the definition of $\tilde{\eth}_{s,p}$, we will have that the sum of the entries of $\eth_{s,p}(x)$ equals $m/m=1$. Consequently, the sum of the entries of $\tilde{\eth}_{s,p}(x)$ equals 1. This completes the proof.
From here on, a map $\tilde{\eth}_{s,p}$ will be called a stochastic $p$-embedding.
@@
Given positive integers $n,p$ and $s\in\{0,1\}^{p+1}$. There are an integer $0<\rho_{p}(n)<d_{s,p}(n)$ and a sparse matrix $R_{s,p}(n)\in\mathbb{R}^{\rho_{p}(n)\times d_{s,p}(n)}$ with $d_{s,p}(n)$ nonzero entries, such that $R_{s,p}(n)\>\tilde{\eth}_{s,p}(x)$ is stochastic and has the least number of non-redundant words (monomial terms) for any $x\in\mathbb{R}^{n}$.
@@
Let $n,p$ be positive integers. Given $s\in\{0,1\}^{p+1}$, consider the structured embedding map $\tilde{\eth}_{s,p}:\mathbb{R}^{n}\to\mathbb{R}^{d_{s,p}(n)}$, where $d_{s,p}(n):=s_{p+1}+\sum_{k=1}^{p}s_{k}n^{k}$ corresponds to the total number of distinct tensor monomials up to degree $p$, plus a constant term.
We aim to construct a sparse matrix $R_{s,p}(n)\in\mathbb{R}^{\rho_{p}(n)\times d_{s,p}(n)}$ that maps the embedding $\tilde{\eth}_{s,p}(x)$ to a stochastic vector of reduced dimension, while preserving all non-redundant monomial terms.
Let us start by defining the matrix $R\in\mathbb{R}^{1\times d_{s,p}(n)}$, with $1$ in its $R_{11}$ entry and with all other entries equal to zero.
For each $2\leq j\leq d_{s,p}(n)$, let us consider the indices $j=k_{1}(j)<k_{2}(j)<\cdots<k_{n_{j}}(j)<d_{s,p}(n)$ that correspond to the same monomial in $\tilde{\eth}_{s,p}(x)$, let us define the matrix $R_{0}\in\mathbb{R}^{1\times d_{s,p}(n)}$ with $1$ in its $R_{1k_{l}(j)}$ entries for $1\leq l\leq n_{j}$, and with all other entries equal to zero. Let us now define the augmented matrix
@@
where $R^{\prime}\in\mathbb{R}^{1\times d_{s,p}(n)}$ is the matrix with entry $R^{\prime}_{1d_{s,p}(n)}$ equal to $1$, and with all other entries equal to 0.
Let us set $R_{s,p}(n):=R$. It can be seen by the way $R$ has been constructed, that the operation $R_{s,p}(n)\tilde{\eth}_{s,p}(x)$ assigns each group of replicates in $\tilde{\eth}_{p}(x)$ to a single representative coordinate of $R_{p}(n)\tilde{\eth}_{p}(x)$. This selection is performed by adding over the redundant entries and replacing this sum into one suitable coordinate. Because of this, and as a consequence of Lemma 1, it is clear that $R_{s,p}(n)\tilde{\eth}_{s,p}(x)$ is stochastic. Let us set $\rho_{p}(n)$ as the number of rows of $R$. This completes the proof.
Using the matrix $R_{s,p}(n)$ described in Lemma 2, we define the reduced stochastic embedding $\tilde{\eth}_{s,p,r}$ as
$\tilde{\eth}_{s,p,r}(x):=R_{s,p}(n)\tilde{\eth}_{s,p}(x),$
@@
For formal implementation purposes, the operation determined by the expression $R_{s,p}(n)\tilde{\eth}_{s,p}(x)$ can be equivalently represented by a vector composed of the distinct monomials of degree less than or equal to $p$, each multiplied by the appropriate integer scaling factor that accounts for its multiplicity in the original tensor expansion.
In this context, a structured reservoir computer refers to a model of the form (1), with $\eth_{p}$ denoting either a reduced stochastic embedding $\tilde{\eth}_{s,p,r}$ or a function of its terms, as illustrated in Section §5. The output coupling matrix $W$ in (1) is stochastic and its structure is determined by a relational graph $\mathcal{G}_{\mathcal{S}}=(V_{\mathcal{S}},E_{\mathcal{S}})$, representing observed or expected interactions among dynamic agents from $t$ to $t+1$ within the training time frame $\{0,1,\ldots,T\}$, for some $T>0$.
@@
[…]
then the existence of a $\mathbb{B}_{\mathcal{S}}(m,n)$-structured stochastic output coupling matrix $W$ that satisfies (1), is a consequence of the following theorem.
@@
The coupling matrices $W$ in equation (1) can be approximately identified in the matrix set $\left(\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(m,n)\right)\cap\mathbb{S}_{m,n}(\mathbb{R})$.
@@
Let $\{x(t)\in\mathbb{R}^{n}\mid t=0,1,\ldots,T\}$ and $\{y(t)\in\mathbb{R}^{m}\mid t=0,1,\ldots,T\}$ denote a stochastic vector time series corresponding to the system’s evolution. We begin by applying a function $\eth_{p}$ of a reduced embedding transformation $\tilde{\eth}_{s,p,r}:=R_{s,p}(n)\tilde{\eth}_{s,p}$ to the observed states, where $p$ is a prescribed tensor order and $R_{s,p}(n)$ is the structure-preserving compression map guaranteed by Lemma 2. This yields a compressed representation of each state, which we collect into the input data matrix
$X_{0}(T):=\begin{bmatrix}|&&|\\ \eth_{p}(x(0))&\cdots&\eth_{p}(x(T-1))\\ |&&|\end{bmatrix}$,
while the elements $y(t)$ are stored in
$X_{1}(T):=\begin{bmatrix}|&&|\\ y(0)&\cdots&y(T-1)\\ |&&|\end{bmatrix}$.
To ensure that the identified coupling matrix respects the structural constraints of the system, we consider a structured dictionary $\mathbb{B}_{\mathcal{S}}(m,n)=\{S_{1},\ldots,S_{q}\}\subset\mathbb{R}^{m\times n}$, derived from the graph $\mathcal{G}_{\mathcal{S}}$ that encodes allowable interactions. Each $S_{j}$ acts as a basis component capturing a localized or interpretable mode of coupling. We construct a matrix $X_{0}$ whose columns correspond to the vectorized actions of these basis elements on the embedded data:
$X_{0}:=\begin{bmatrix}\text{vec}(S_{1}X_{0}(T))&\cdots&\text{vec}(S_{q}X_{0}(T))\end{bmatrix}$,
and define the target vector as $X_{1}:=\text{vec}(X_{1}(T))$.
The identification task is then posed as the following structured nonnegative least squares problem:
$\mathbf{a}:=\arg\min_{\hat{\mathbf{a}}\in(\mathbb{R}_{0}^{+})^{q}}\left\|\begin{bmatrix}X_{0}^{\top}X_{0}\\ C\end{bmatrix}\hat{\mathbf{a}}-\begin{bmatrix}X_{0}^{\top}X_{1}\\ \mathbf{1}_{p}\end{bmatrix}\right\|_{2},$
where $\|\cdot\|_{2}$ denotes the Euclidean vector norm, and $C$ is a constraint matrix chosen to enforce the membership of the resulting linear combination $\hat{W}:=\sum_{j=1}^{q}\mathbf{a}[j]S_{j}$ in the structured subset $\mathbb{S}_{m,n}(\mathbb{R})$. Such constraints may include nonnegativity, block sparsity, or stochasticity properties, depending on the system’s prior assumptions.
As discussed in BOUTSIDIS2009760, the above system corresponds to a convex quadratic optimization problem with linear constraints (e.g., a Nonnegative Least Squares problem), and is therefore solvable in polynomial time up to arbitrary precision.
Since $\hat{W}$ is constructed as a linear combination of elements in $\mathbb{B}_{\mathcal{S}}(m,n)$, it lies in $\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(m,n)$ by definition. The enforcement of structural constraints via $C$ ensures that $\hat{W}\in\mathbb{S}_{m,n}(\mathbb{R})$. Hence,
$\hat{W}\in\left(\mathrm{span}\>\mathbb{B}_{\mathcal{S}}(m,n)\right)\cap\mathbb{S}_{m,n}(\mathbb{R}),$